# Scaling-aware rating of count forecasts

### v3

@@ Appendix A Heuristic algorithm for non-parametric Expectation Maximization
For the purpose of discussing our rating methodology on “perfect” Poissonian forecasts for a given dataset of sales values $s_{j}$, we construct predictions using the following heuristic, which is equivalent to Expectation Maximization: Given the set of observed sales $s_{j}$, we assume that these integer outcomes are the results of Poisson processes with rates $t_{j}$, i.e. that Eq. (4) holds. Given these known observations $s_{j}$, we need to evaluate $P(t|s)$, i.e. the conditional probability to have predicted $t$ given that we do observe $s$, to create plausible predictions $t_{j}$ (Tichy, 2023b). Remember that Eq. (4) states the conditional probability to observe $s$, given that $t$ is predicted. Bayes’ rule gives
@@
To provide a set of plausible predictions $t_{j}$, we sample one prediction $t_{j}$ for each observation $s_{j}$ via $t_{j}\sim P(t|s_{j})$. Applying the iterative procedure on all data categories (product groups, locations…) separately, we obtain unbiased Poissonian forecasts that could have yielded the set of sales $s_{j}$. We have experienced that a dozen iterations is typically sufficient to reach a good artificial forecast, i.e. a forecast for which $P^{(k)}_{\text{observation}}(s)\approx P_{\text{observation}}(s)$ while, for each prediction $t$, the resulting sales $s$ are Poisson-distributed, as shown in Figure 1.
A.1 Treatment of predicted 0-observations
In the update rule (16), the cases $P^{(k)}_{\text{observation}}(s)=0$ and/or $P_{\text{observation}}(s)=0$ need special attention.
@@ A.1 Treatment of predicted 0-observations
    When $P_{\text{observation}}^{(k)}(s)=P_{\text{observation}}(s)=0$, we set the quotient ${P_{\text{observation}}(0)}/{P^{(k)}_{\text{observation}}(0)}$ to 1: The probability 0 for the observation $s$ predicted by $P_{\text{observation}}^{(k)}(s)$ matches the empirical frequency $P_{\text{observation}}(s)$, no adjustment is needed.
@@
    When $P_{\text{observation}}(s)>0$ but $P_{\text{observation}}^{(k)}(s)=0$, we need to “boost” components that contribute to that observation, hence we set the otherwise undefined quotient to 2.

### v4

@@ Appendix A Heuristic algorithm for non-parametric Expectation Maximization
For the purpose of discussing our rating methodology on “perfect” Poissonian forecasts for a given dataset of sales values $s_{j}$, we construct predictions using the following heuristic, which is equivalent to Expectation Maximization: Given the set of observed sales $s_{j}$, we assume that these integer outcomes are the results of Poisson processes with rates $t_{j}$, i.e. that Eq. (4) holds. Given the known observations $s_{j}$, we need to evaluate $P(t|s)$, i.e. the conditional probability to have predicted $t$ given that we do observe $s$, to create plausible predictions $t_{j}$ (Tichy, 2023b). Remember that Eq. (4) states the conditional probability to observe $s$, given that $t$ is predicted. Bayes’ rule gives
@@
To provide a set of plausible predictions $t_{j}$, we sample one prediction $t_{j}$ for each observation $s_{j}$ via $t_{j}\sim P(t|s_{j})$. Applying the iterative procedure on all data categories (product groups, locations…) separately, we obtain unbiased Poissonian forecasts that could have yielded the set of sales $s_{j}$. We have experienced that a dozen iterations is typically sufficient to reach a good artificial forecast, i.e. a forecast for which $P^{(k)}_{\text{observation}}(s)\approx P_{\text{observation}}(s)$ while, for each prediction $t$, the resulting sales $s$ are Poisson-distributed, as shown in Figure 1. In practice, we discretize the space of rates $t$ into a granular array of length $G=5^{\prime}000$, $t_{0}\dots t_{G}$, such that (16) is performed on all $P_{k}(t_{0})\dots P_{k}(t_{G})$ in a numerical fashion.
A.1 Treatment of improbable observations
In the update rule (16), the cases $P^{(k)}_{\text{observation}}(s)=0$ and/or $P_{\text{observation}}(s)=0$ need special attention, especially in practice, when one recurs to numerical discretization of the space of rates $t$.
@@ A.1 Treatment of improbable observations
    When $P_{\text{observation}}^{(k)}(s)=P_{\text{observation}}(s)=0$, we set the quotient of ${P_{\text{observation}}(0)}$ and ${P^{(k)}_{\text{observation}}(0)}$ to 1: The probability 0 for the observation $s$ predicted by $P_{\text{observation}}^{(k)}(s)$ matches the empirical frequency $P_{\text{observation}}(s)$, no adjustment is needed.
@@
    When $P_{\text{observation}}(s)>0$ but $P_{\text{observation}}^{(k)}(s)=0$ (or very close to 0), we need to “boost” components that contribute to that observation $s$, hence we set the otherwise undefined (or very large) quotient to 2.